OpenStudy (anonymous):

What is the length of the diagonal of a rectangle whose width is 130 centimeters and whose length is 200 centimeters? (Round your answer to the nearest whole centimeter.) 229 cm        234 cm        237 cm        239 cm

OpenStudy (anonymous):

Use the property of a right triangle \[ d^2 = w^2+ l^2 \]

OpenStudy (anonymous):

\[ d^2 = 200^2 + 130^2 \]

OpenStudy (anonymous):

It should be easy now.

OpenStudy (anonymous):

The angles of a rectangle are each of value 90 deg. So , you can proceed as eliassaab showed and find out the value of the diagonal from it.

OpenStudy (anonymous):

\[ d^2 = 200^2 + 130^2=56900\\ d= \sqrt{56900}\approx 238.537 \]

OpenStudy (anonymous):

Thanks, lol I called myself trying to figure it out.